Projectile Motion: Other applications


A boy throws a rock with speed y = 18.3 m/s at an angle of ∅ = 57.0
1
Expert's answer
2021-09-28T11:13:44-0400

The range of the rock is


"R=\frac{v^2\sin[2\alpha]}{g}=\frac{18.3^2\sin[2·57°]}{9.8}=31.2\text m."